# Towards Inclusive Face Recognition Through Synthetic Ethnicity Alteration thanks: This work was supported by the European Union’s Horizon 2020 Research and Innovation Program under Grant 883356

### v1

@@ I Introduction
Automated face recognition has made significant strides due to the rapid advancement of deep learning algorithms. The current state-of-the-art FRS [1, 2] exhibits exceptionally high accuracy across different face recognition benchmarks in unconstrained environments (e.g., LFW [3], IJB-A [4], AgeDB [5], ICD [6], and Mega Asia [7]), with performance almost reaching near-ideal levels. Several benchmarks focus on specific challenges, such as pose and aging [8], which lead to substantial performance drops on other facial covariates. Despite improved performance, various studies have however questioned if face recognition systems perform equally well across various demographic groups [9]. Many FRS have exhibited lower accuracy rates for certain demographic groups with established empirical evidence leading to strong critique on biased performance [9] with commercial systems not being an exception [^1] https://news.mit.edu/2018/study-finds-gender-skin-type-bias-artificial-intelligence-systems-0212. Not different from FRS is the recognition phenomenon in humans who tend to identify/recognize people of one’s own race, often referred to as the other-race effect [10, 11].
@@
Synthetic data has emerged as a promising alternative to real-world data, offering greater control over demographics and facial attributes manipulation. This allows for alignment with the desired demographic distributions while maintaining consistent verification and face image quality across diverse demographic groups [17]. Moreover, synthetic face generators possess the potential not only to produce virtually infinite data, but also to mitigate privacy concerns that have prompted the discontinuation of established datasets and the enactment of regulatory frameworks such as the EU-GDPR [18]. We motivate our work taking the advantage of synthetic data generation to create identity of same person across different ethnicity to study the suitability for diversifying the datasets for training FRS. Further, some individuals may exhibit features of different ethnicities before and after due to extensive cosmetic surgeries [19] or skin disorder disease [20] as shown in Fig. 2. A significant transformation in facial features around the nose and skin tone can be noticed from Fig. 2. We thus assert transforming the ethnicity of a person across different ethnicities can help in diversifying and increasing the datasets for training FRS as shown in Fig 1. To validate our assertion, we pose three critical questions: Q1. Does synthetic ethnicity alteration and skin tone modification using Generative Adversarial Networks (GANs) find use in FRS for representing a subject in different ethnicities? Q2. Does synthetic ethnicity alteration and skin tone modification of a subject’s facial image resemble ITA of real faces? Q3. Do FIQA and FRS measures for synthetic ethnicity-altered image datasets indicate dependable performance, similar to real datasets of different ethnicities?
@@
In the course of answering this question, we first construct a balanced dataset of multiple ethnicities (i.e., Indian, Asian and Black) and make use of it in generating transformed ethnicities. We specifically make use of existing image-to-image translation models and manifold learning based translation models to generate different ethnicity representation of any subject. We generate a new synthetic dataset from each of these settings resulting in $45,000$ images representing three ethnicities. We then systematically assess the skin-tone of transformed ethnicity to verify the representative nature of skin-tone of real data using ITA. We then assess if the images provide performances similar to real data by analysing the FIQA and subsequently FRS. For the analysis using FIQA, we provide Error-vs-Discard Curves (EDC) using three different FRS. Complementary to that, we use another four FRS to benchmark the verification performance in a cross-ethnicity scenario (e.g., a subject from Indian ethnicity represented in Asian). Our contributions in this work include:
@@ III Investigation Methodology
The primary objective of this study was to study the feasibility of altering the ethnicity synthetically for FRS applications. To accomplish this, a well-structured and balanced dataset which we refer to as Ethnicity Alteration Training (ETAT) dataset (detailed in Section IV) was created representing three distinct ethnic groups (Asian, Black, and Indian). The GAN-based models were trained to alter the ethnicities using six different combinations: Asian $\mapsto$ Black, Asian $\mapsto$ Indian, Black $\mapsto$ Asian , Black $\mapsto$ Indian , Indian $\mapsto$ Asian, and Indian $\mapsto$ Black. The trained models were applied to the RFW [12] dataset to perform the ethnicity transformations. We first analyze the skin-tones of generated images. Further to that, FIQA using Error-vs-Discard characteristic (EDC) analysis is conducted to measure the suitability of ethnicity-altered images. To assess the effects of ethnicity alterations on FRS, four pre-trained deep FRS models (FaceNet, PFE, Elastic Arc+, and Elastic Cos+) were studied. The results were measured using three complementary metrics: (i) cross-ethnicity skin tone transformation utilizing an ITA, (ii) FIQA analysis using EDC and (iii)cross-ethnicity face verification performance using FNMR and FMR.
@@ IV-A Ethnicity Alteration Training (ETAT) Dataset
The process of developing the ETAT dataset involved initially predicting the race of images within existing databases and then adding images of Indian ethnicity through dedicated web crawling efforts [^2] https://github.com/deepanprabhu/duckduckgo-images-api. This subset of web-crawled images constitutes a comprehensive Indian dataset from various states [^3] The detailed distribution is shown in the supplementary material due to page constraints. containing approximately 57,000 images exclusively from Indians. For ground truth annotation in publicly available datasets (e.g., CelebA [31], FFHQ [30], AgeDB [5], Mega Asia [34], and CASIA-WebFace [35]), state-of-the-art methods (e.g., DeepFace [36]) were utilized to predict ethnicity and age in these datasets. All images were processed using an advanced face detection technique (i.e., MTCNN [37]), including cropping, alignment, and resizing ($128\times 128$), to ensure consistency and quality. After filtering out blurry, non-readable, and low-quality images, $15,000$ high-quality images representing Asian, Black, and Indian ethnicities were selected for training. This is the first Indian subset to represent an Indian dataset from different states (detail is provided in supplementary material). The ETAT dataset fills a crucial gap in Indian face representation in the FRS and facilitates research on mitigating race bias for more inclusive AI systems.
@@ VI Quantitative Evaluation
    Face quality assessment performance: This assessment presents an in-depth analysis of the correlation between FRS and FIQA on an altered image of a different ethnicity [39]. To evaluate the face quality assessment performance, we follow the methodology by Torsten et al. [40] using Error versus discard characteristics (EDC).
@@ VI-A Skin Tone Estimation
TABLE I: Cross ethnicity face verification rate (%) of face recognition systems on Asian (Genuine:11,443, Imposter:14305), Black (Genuine:10,195, Imposter:15,036), and Indian (Genuine:10,143, Imposter:13,097) ethnicity. The FNMR is evaluated at two FMR thresholds for four face recognition systems.
@@ VI-B Face image quality assessment
We further evaluate the utility of synthetically altered images by studying the EDC to verify the image quality usable within FRS. EDC curve shows the verification error rate on the y-axis achieved when discarding a certain percentage of face images (x-axis). Based on the predicted quality values, these discarded images have the lowest predicted quality, and the error rate is calculated on the remaining images. EDC curve indicates good quality estimation when the verification error decreases consistently when the ratio of unconsidered images increases. We employ the Arcface FRS described to measure the errors vs discard characteristics. The verification error is reported regarding the false non-match rate (FNMR) at fixed false match rates (FMR). The FMR is reported at 0.05% FMR threshold.
@@ VI-C Cross Ethnicity Face Verification
We thoroughly evaluated the FRS on the RFW test datasets by creating genuine and impostor image pairs. To create genuine pairs, we utilized the generated images of each target domain and the input images of the source domain ethnicity for individual subjects. Furthermore, we randomly generated cross-subject imposter pairs for comparison. The results of SOTA FRS evaluation are presented in Table I and illustrate the system’s performance at the FNMR with an FMR of $0.1\%$ and $0.01\%$.
It can be noticed that CycleGAN and StarGAN verification rates are lower than FGAN due to the latter generating better ethnicity-altered images (as seen in Figure 3, Figure 4, and Figure 5) and the distribution of ITA angles between altered and real images overlaps, as illustrated in Figure 7. Also, altered Asian and Black images using CycleGAN exhibit minor differences in verification accuracy compared to real and cross-ethnicity altered images. Additionally, altered Indian ethnicity images using StarGAN demonstrate better verification accuracy than those altered using CycleGAN, suggesting that these models are more effective for generating cross-altered ethnicity synthetic datasets. On an average, all FRS perform better for Indian ethnicity than Asian and Black ethnicity across real and altered images. Indeed, it is notable that the accuracy of cross-ethnicity face verification increases when utilizing newer models such as Elastic-Arc+ and Elastic-Cos+, as opposed to older models such as FaceNet and PFE, both for genuine and imposter score distributions on both real and altered images. The supplementary material provides the genuine and imposter score distributions of the top-performing ElasticCos+ results.

### v2

@@ I Introduction
Automated face recognition has made significant strides due to the rapid advancement of deep learning algorithms. The current state-of-the-art FRS [1, 2] exhibits exceptionally high accuracy across different face recognition benchmarks in unconstrained environments (e.g., LFW [3], IJB-A [4], AgeDB [5], ICD [6], and Mega Asia [7]), with performance almost reaching near-ideal levels. Several benchmarks focus on specific challenges, such as pose and aging [8], which lead to substantial performance drops on other facial covariates. Despite improved performance, various studies have however questioned if face recognition systems perform equally well across various demographic groups [9]. Many FRS have exhibited lower accuracy rates for certain demographic groups with established empirical evidence leading to strong critique on biased performance [9] with commercial systems not being an exception [^1] https://news.mit.edu/2018/study-finds-gender-skin-type-bias-artificial-intelligence-systems-0212. Not different from FRS is the recognition phenomenon in humans who tend identify/recognize people of one’s own race, often referred to as the other-race effect [10, 11].
@@
Synthetic data has emerged as a promising alternative to real-world data, offering greater control over demographics and facial attributes manipulation. This allows for alignment with the desired demographic distributions while maintaining consistent verification and face image quality across diverse demographic groups [17]. Moreover, synthetic face generators possess the potential not only to produce virtually infinite data, but also to mitigate privacy concerns that have prompted the discontinuation of established datasets and the enactment of regulatory frameworks such as the EU-GDPR [18]. We motivate our work taking the advantage of synthetic data generation to create identity of same person across different ethnicity to study the suitability for diversifying the datasets for training FRS. Further, some individuals may exhibit features of different ethnicities before and after due to extensive cosmetic surgeries [19] or skin disorder disease [20] as shown in Fig. 2. A significant transformation in facial features around the nose and skin tone can be noticed from Fig. 2. We thus assert transforming the ethnicity of a person across different ethnicities can help in diversifying and increasing the datasets for training FRS as shown in Fig 1. To validate our assertion, we pose three research questions: Q1. Does synthetic ethnicity alteration and skin tone modification using Generative Adversarial Networks (GANs) find use in FRS for representing a subject in different ethnicities? Q2. Does synthetic ethnicity alteration and skin tone modification of a subject’s facial image resemble ITA of real faces? Q3. Do FIQA and FRS measures for synthetic ethnicity-altered image datasets indicate dependable performance, similar to real datasets of different ethnicities?
@@
In the course of answering these questions, we first construct a balanced dataset of multiple ethnicities (i.e., Indian, Asian and Black) and make use of generating transformed ethnicities. We specifically make use of existing image-to-image translation models and manifold learning based translation models to generate different ethnicity representation of any subject. We generate a new synthetic dataset from each of these settings resulting in $45,000$ images representing three ethnicities. We then systematically assess the skin-tone of transformed ethnicity to verify the representative nature of skin-tone of real data using ITA. We then assess if the images provide performances similar to real data by analysing the FIQA and subsequently FRS. For the analysis using FIQA, we provide Error-vs-Discard Characteristic curves (EDC) using three different FRS. Complementary to that, we use another fourth FRS to benchmark the verification performance in a cross-ethnicity scenario (e.g., a subject from Indian ethnicity represented in Asian). Our contributions in this work include:
@@ III Investigation Methodology
The primary objective of this work was to study the feasibility of altering the ethnicity synthetically for FRS applications. To accomplish this, a well-structured and balanced dataset which we refer to as Ethnicity Alteration Training (ETAT) dataset (detailed in Section IV) was created representing three distinct ethnic groups (Asian, Black, and Indian). The GAN-based models were trained to alter the ethnicities using six different combinations: Asian $\mapsto$ Black, Asian $\mapsto$ Indian, Black $\mapsto$ Asian , Black $\mapsto$ Indian , Indian $\mapsto$ Asian, and Indian $\mapsto$ Black. The trained models were applied to the RFW [12] dataset to perform the ethnicity transformations. We first analyze the skin-tones of generated images. Further to that, FIQA using Error-vs-Discard characteristic (EDC)  [^2] https://www.iso.org/standard/79519.html analysis is conducted to measure the suitability of ethnicity-altered images. To assess the effects of ethnicity alterations on FRS, four pre-trained deep FRS models (FaceNet, PFE, Elastic Arc+, and Elastic Cos+) were studied. The results were measured using three complementary metrics: (i) cross-ethnicity skin tone transformation utilizing an ITA, (ii) FIQA analysis using EDC and (iii) cross-ethnicity face verification performance using FNMR and FMR.
@@ IV-A Ethnicity Alteration Training (ETAT) Dataset
The process of developing the ETAT dataset involved initially predicting the race of images within existing databases and then adding images of Indian ethnicity through dedicated web crawling efforts [^3] https://github.com/deepanprabhu/duckduckgo-images-api. This subset of web-crawled images constitutes a comprehensive Indian dataset from various states containing approximately 57,000 images exclusively from Indians. For ground truth annotation in publicly available datasets (e.g., CelebA [31], FFHQ [30], AgeDB [5], Mega Asia [34], and CASIA-WebFace [35]), state-of-the-art methods (e.g., DeepFace [36]) were utilized to predict ethnicity and age in these datasets. All images were processed using an advanced face detection technique (i.e., MTCNN [37]), including cropping, alignment, and resizing ($128\times 128$), to ensure consistency and quality. After filtering out blurry, non-readable, and low-quality images, $15,000$ high-quality images representing Asian, Black, and Indian ethnicities were selected for training. This is the first Indian subset to represent an Indian dataset from different states (details are provided in supplementary material). The ETAT dataset fills a crucial gap in Indian face representation in the FRS and facilitates research on mitigating race bias for more inclusive AI systems.
@@ VI Quantitative Evaluation
    Face image quality assessment performance: This assessment presents an in-depth analysis of the correlation between FRS and FIQA on an altered image of a different ethnicity [39]. To evaluate the face image quality assessment performance, we follow the methodology by Schlett et al. [40] using Error versus discard characteristics (EDC).
@@ VI-A Skin Tone Estimation
TABLE I: Cross ethnicity face verification rate (%) of face recognition systems on Asian (mated comparison:11,443, non-mated comparison:14305), Black (mated comparison:10,195, non-mated comparison:15,036), and Indian (mated comparison:10,143, non-mated comparison:13,097) ethnicity. The FNMR is evaluated at two FMR thresholds for four face recognition systems.
@@ VI-B Face image quality assessment
We further evaluate the utility of synthetically altered images by studying the EDC to verify the image quality usable within FRS. EDC curve shows the false non-match error rate on the y-axis achieved when discarding a certain percentage of low-quality images (x-axis). Based on the predicted quality values, these discarded images have the lowest predicted quality, and the error rate is calculated on the remaining images. EDC curve indicates good quality estimation when the verification error decreases consistently when the ratio of unconsidered images increases. We employ the Arcface FRS described to measure the errors vs discard characteristics. The verification error is reported regarding the false non-match rate (FNMR) at fixed false match rates (FMR). The FMR is reported at 0.05% FMR threshold.
@@ VI-C Cross Ethnicity Face Verification
We thoroughly evaluated the FRS on the RFW test datasets by creating mated and non-mated image pairs. To create mated pairs, we utilized the generated images of each target domain and the input images of the source domain ethnicity for individual subjects. Furthermore, we randomly generated cross-subject imposter pairs for comparison. The results of SOTA FRS evaluation are presented in Table I and illustrate the system’s performance at the FNMR with an FMR of $0.1\%$ and $0.01\%$.
It can be noticed that CycleGAN and StarGAN verification rates are lower than FGAN due to the latter generating better ethnicity-altered images (as seen in Figure 3, Figure 4, and Figure 5) and the distribution of ITA angles between altered and real images overlaps, as illustrated in Figure 7. Also, altered Asian and Black images using CycleGAN exhibit minor differences in verification accuracy compared to real and cross-ethnicity altered images. Additionally, altered Indian ethnicity images using StarGAN demonstrate better verification accuracy than those altered using CycleGAN, suggesting that these models are more effective for generating cross-altered ethnicity synthetic datasets. On an average, all FRS perform better for Indian ethnicity than Asian and Black ethnicity across real and altered images. Indeed, it is notable that the accuracy of cross-ethnicity face verification increases when utilizing newer models such as Elastic-Arc+ and Elastic-Cos+, as opposed to older models such as FaceNet and PFE, both for mated and non-mated score distributions on both real and altered images.